When a child breaks open her piggy bank, she finds a total of 64 coins, consisting of nickels, dimes, and quarters. The total value of the coins is $6. If the nickels were dimes, and the dimes were nickels, the value of the coins would be $5. How many nickels, dimes, and quarters were in the piggy bank?
Answer by richwmiller(17219)

n+d+q=64,
5n+10d+25q=600,
10n+5d+25q=500
d = 40, n = 20, q = 4